OpenStudy (anonymous):

Based on the table of values below, find the slope between points where x = 5 and where x = 10. x y 5 2 7 6 10 7 −1 one half 1 2

OpenStudy (anonymous):

any ideas

OpenStudy (anonymous):



OpenStudy (texaschic101):

slope formula : (y2 - y1) / (x2 - x1) (5,2) x1 = 5 and y1 = 2 (10,7) x2 = 10 and y2 = 7 now we sub slope = (7 - 2) / (10 - 5) slope = 5/5 = 1

OpenStudy (anonymous):

thank you it helps

OpenStudy (texaschic101):

your quite welcome...always glad to help :)